Middle C on a piano has a frequency of 256 Hz. If the speed of sound in air is 343 m/s, what is the wave’s period?
1
Expert's answer
2020-11-27T14:07:29-0500

We can find the period of the wave from the formula:


"f=\dfrac{1}{T},""T=\dfrac{1}{f}=\dfrac{1}{256\ Hz}=3.9\cdot10^{-3}\ s."

Answer:

"T=3.9\cdot10^{-3}\ s."